OpenStudy (haleyelizabeth2017):

What are the real and complex solutions of the polynomial equation?

OpenStudy (haleyelizabeth2017):

\(27x^3+125=0\) I got x=-5/3 so far, I just am confused on what to do next....

OpenStudy (mathstudent55):

Did you factor the binomial? Do you know which kind of factoring you need?

OpenStudy (haleyelizabeth2017):

huh? Don't I need to divide using the synthetic division?

OpenStudy (mathstudent55):

Hint: \(27x^3 + 125 = (3x)^3 + (5)^3\)

OpenStudy (haleyelizabeth2017):

I got the x=-5/3 by doing this: \(27x^3=-125\) -->subtract 125 both sides \(x^3=-125/27\) -->divide both sides by 27 \(x=-5/3\) --> cube root both -125 and 27 .......is this correct?

OpenStudy (mathstudent55):

This is not how you solve this kind of problem. You need to factor the left side. Then set each factor equal to zero. You will get three solutions.

OpenStudy (mathstudent55):

The binomial you have is a sum of cubes. There is a special kind of factoring for a sum of cubes. You may be familiar with a perfect square trinomial, and with the difference of two squares which are all special cases of factoring. This is another one of those special cases.

OpenStudy (mathstudent55):

Here is how a sum of two cubes is factored: \(a^3 + b^3 = (a + b)(a^2 - ab + b^2)\)

OpenStudy (mathstudent55):

In your case, a = 3x, and b = 5.

OpenStudy (haleyelizabeth2017):

okay

OpenStudy (haleyelizabeth2017):

I am like really confused now......:(

OpenStudy (mathstudent55):

Use the factoring formula I gave you ,and use 3x for a and 5 for b. You should get this: \((3x)^3 + (5)^3 = (3x + 5)[(3x)^2 - (3x)(5) + (5)^2] \)

OpenStudy (haleyelizabeth2017):

okay

OpenStudy (mathstudent55):

Now you need to simplify what is inside the square brackets.

OpenStudy (haleyelizabeth2017):

okay so 3x^2-15x+25?

OpenStudy (haleyelizabeth2017):

\((3x+5)(3x^2-15x+25)\)

OpenStudy (mathstudent55):

Almost. Notice the first term is (3x)^2. The entire 3x quantity has to be squared. That means the 3 is also squared.

OpenStudy (mathstudent55):

I mean in the second parentheses.

OpenStudy (haleyelizabeth2017):

okay

OpenStudy (mathstudent55):

\((3x+5)(\color{red}{9}x^2-15x+25) = 0\)

OpenStudy (mathstudent55):

Got it?

OpenStudy (haleyelizabeth2017):

not really........:( sorry I'm normally good at math but this is just confusing :'(

OpenStudy (mathstudent55):

You are very close. Look at the 9 in red above.

OpenStudy (haleyelizabeth2017):

okay

OpenStudy (mathstudent55):

In the square brackets, the first term was (3x)^2 That means the quantity 3x must be squared, so (3x)^2 = 3x * 3x = 9x^2

OpenStudy (haleyelizabeth2017):

oh!

OpenStudy (haleyelizabeth2017):

now I get it :D

OpenStudy (mathstudent55):

Ok, now we have correctly factored it. Now we set each factor equal to zero, and we solve each equation.

OpenStudy (haleyelizabeth2017):

how do we do the \(9x^2-15x+25\)=0?

OpenStudy (haleyelizabeth2017):

do we have to do that one by completing the square?

OpenStudy (mathstudent55):

\((3x+5)(9x^2-15x+25)=0\) \(3x+5 = 0 ~~~or~~~ 9x^2-15x+25=0\) The left one is easy, and you know how to solve it. The right one, you need th equadratic formula.

OpenStudy (haleyelizabeth2017):

oh! I get it now lol :P thanks one sec

OpenStudy (mathstudent55):

Yes, completing the square will also work, but I think the quadratic formula is easier.

OpenStudy (mathstudent55):

\(x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a} \)

OpenStudy (haleyelizabeth2017):

okay.....

OpenStudy (mathstudent55):

For the quadratic formula, a = 9, b = -15, c = 25

OpenStudy (haleyelizabeth2017):

\(x = \dfrac{15 \pm \sqrt{-15^2 - 4(9)(25)}}{2(9)}\)

OpenStudy (haleyelizabeth2017):

\(x = \dfrac{15 \pm \sqrt{ 225- 900}}{18}\)

OpenStudy (mathstudent55):

Almost. \(x = \dfrac{15 \pm \sqrt{(-15)^2 - 4(9)(25)}}{2(9)}\)

OpenStudy (haleyelizabeth2017):

what's different?

OpenStudy (mathstudent55):

You calculated it correctly, but you need the parentheses around the -15 when you square -15.

OpenStudy (haleyelizabeth2017):

oh yeah.....haha whoops

OpenStudy (mathstudent55):

\(-15^2 = -225\) This is the opposite of 15 squared. \((-15)^2 = 225\) This is the number -15 squared.

OpenStudy (haleyelizabeth2017):

yep....I get my mistake now lol whoopsies so now: \(x = \dfrac{15 \pm \sqrt{-675}}{18}\)

OpenStudy (haleyelizabeth2017):

how can I simplify that now?

OpenStudy (mathstudent55):

\(x = \dfrac{15 \pm \sqrt{(-675)}}{18}\) Yes, I got the same.

OpenStudy (mathstudent55):

675 = 3 * 225, and 225 = 15^2

OpenStudy (haleyelizabeth2017):

so would you just put \(3(15)^2\) in sqrt?

OpenStudy (haleyelizabeth2017):

I don't know that part.....

OpenStudy (mathstudent55):

Yes, but remember it is -675, so you need to take out i.

OpenStudy (haleyelizabeth2017):

\[\sqrt{3(15)^2}*i\]

OpenStudy (mathstudent55):

Yes, you are correct. \(x = \dfrac{15 \pm i\sqrt{-675}}{18}\) \(x = \dfrac{15 \pm i\sqrt{675}}{18}\) \(x = \dfrac{15 \pm i\sqrt{3(15)^2}}{18}\) \(x = \dfrac{15 \pm 15i\sqrt{3}}{18}\)

OpenStudy (haleyelizabeth2017):

okay. is that the answer?

OpenStudy (mathstudent55):

Sorry, I really have to go, so I'll just finish it. Please continue working and use my solution to check your work.

OpenStudy (haleyelizabeth2017):

so \(x = \dfrac{15 \pm 15i\sqrt{3}}{18}\) and x=-5/3

OpenStudy (haleyelizabeth2017):

or is it "or"

OpenStudy (mathstudent55):

15 and 18 are both divisible by 3, so we can reduce the fraction.

OpenStudy (haleyelizabeth2017):

okay thanks! is that all I have to do?

OpenStudy (haleyelizabeth2017):

x = \(\dfrac{5 \pm 15i\sqrt{3}}{6}\)

OpenStudy (mathstudent55):

\(x = \dfrac{15 \pm 15i\sqrt{3}}{18}\) \(x = \dfrac{5 \pm 5i\sqrt{3}}{6}\)

OpenStudy (mathstudent55):

Almost. The second 15 in the numerator also needs to be divided by 3.

OpenStudy (haleyelizabeth2017):

oh! okay thank you so much for all of your help! I get it now........

OpenStudy (haleyelizabeth2017):

:D

OpenStudy (mathstudent55):

Then you can separate the plus and the minus and write as two solutions and also remember the simple solution you already had, the real one. All three solutions are separated by the word "or."

OpenStudy (haleyelizabeth2017):

okay :) thanks again!

OpenStudy (mathstudent55):

\(x = -\dfrac{5}{3} \) or \(x = \dfrac{5 + 5i\sqrt{3}}{6}\) or \(x = \dfrac{5 - 5i\sqrt{3}}{6}\)

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (haleyelizabeth2017):

:D